# Effective Quantization of Muon Optimizer States

### v1

@@ 4.3 The 8-bit Muon algorithm
Consider a step of Muon with momentum (using the exact polar factor rather than the Newton–Schulz approximation) with and without 8-bit linear quantization of the momentum. Let the layer weights and momentum state be $\mathbf{W}^{(t-1)}$ and $\mathbf{M}^{(t-1)}$, each with $d$ entries (see Appendix C.4 for full update formulas).
@@
5 Experiments
5.1 Pre-training with 8-bit Muon
Architectures For the pre-training task, we train from scratch a modified version of the GPT2 architecture (Radford et al., 2019) in which the learned positional embeddings are replaced by rotary positional embeddings (RoPE) (Su et al., 2024). To understand the scaling effect, we consider 4 different sizes - XS (97M), Small (124M), Medium (405M) and XL (1.6B). For the sake of brevity, we will refer to this architecture as GPT in the rest of the paper. Detailed notes on the architecture can be found in Table 6 in Appendix A.1.2.
Datasets Our pre-training dataset consists of 2 billion tokens from the FineWeb-Edu dataset Penedo et al. (2024) for the XS, Small and Medium models, and 4 billion tokens for the XL model. To achieve this, we take twenty (forty for the XL variant) chunks of the dataset, each chunk being approximately 100M tokens. We use 50k samples from the validation split of FineWeb to measure validation loss, totaling approximately 100M tokens.
Training details We pre-train the 4 GPT models from scratch using the splits from FineWeb described earlier. For AdamW, we set $\beta_{1}$ and $\beta_{2}$ to 0.9 and 0.999 respectively. The $\epsilon$ value for AdamW variants is set to $10^{-8}$. For Muon, we set the momentum parameter to 0.95.
The block size of the quantized versions of AdamW and Muon is set to 2048. We use the bitsandbytes library (Dettmers et al., 2023) for quantization / dequantization routines for linear and dynamic quantization. We fix decoupled weight decay to 0.1 for all experiments.
We use the WSD learning rate schedule (Hu et al., 2024), with a linear warm up of the learning rate from 0 to the peak in 10% of training steps, a linear decay to zero in the last 10% steps. For XS, Small and Medium models, we fix the peak learning rate to $3\times 10^{-4}$ for all variants. For the XL model, we reduced the maximum learning rate to $2\times 10^{-4}$ for all variants. Since most of the peak learning rates have been tuned for AdamW, we make no attempt to tune the peak learning rate for Muon and its variants, since our version of Muon is supposed to be a drop-in replacement for AdamW without any requirement to tune learning rate or weight decay (Liu et al., 2025).
We use distributed data parallel (DDP) for all experiments, with training working on multiple GPUs at the same. Our largest model is 1.6B and fits on a single NVIDIA H100 GPU. The global batch size is fixed at 262k tokens, and is achieved with varying gradient accumulation across model sizes. A detailed table on hyperparameters and training configurations can be found in Appendix A.1.2.
Evaluation criteria We use validation loss as the primary criterion for comparing performance for pre-training. We exhibit results from a single representative run of each experimental variant since the variance across runs was low.
5.1.1 Results
111.251.251.51.51.751.7522334455Tokens (billions)Validation lossMuon-8D/Adam-32Muon-8L/Adam-32Muon-8DMuon-8L
Figure 2: Training curves on GPT-Medium comparing four variants of 8-bit Muon (zoomed in version, from 1B to 2B tokens). Except Muon-8L, all the others follow each other closely.
Investigating the stability of 8-bit Muon optimizers In Figure 2, we show the results of comparing all variants of 8-bit Muon described in Table 1 on the task of pre-training the GPT-Medium model. Before testing these variants, we confirmed that AdamW-8L diverged on this problem early during training. Interestingly, Muon-8L did not diverge, but performed poorly. Remarkably, all other versions of 8-bit Muon closely match each other. This includes Muon-8L/Adam-32, a version that uses linear quantization for Muon-associated parameters. This corroborates our theoretical findings that AdamW is particularly unstable especially when using simple techniques like linear quantization. Because of these findings, we avoid the Muon-8L variant and instead use alternatives like Muon-8L/Adam-32.
The main results can be found in Figure 3 and Table 3. We make the following observations:
Muon outperforms AdamW variants Unsurprisingly, Muon outperforms all variants of AdamW. This is despite no tuning of the peak learning rate or weight decay.
8-bit Muon variants are competitive with Muon Surprisingly, all the 8-bit variants of Muon achieve a loss within a $\sim 1\%-2\%$ error margin of Muon for XS/Small/Medium, comfortably outperforming AdamW-32 and Adam-8D. This includes Muon-8L/Adam-32, which uses linear quantization for hidden matrix-valued parameters. For the XL model, the 8-bit Muon variants match the performance of Muon. These results show that Muon is robust to the underlying quantization scheme for state compression.
00.50.5111.51.522445566Validation lossGPT XS00.50.5111.51.52233445566GPT Small00.50.5111.51.52233445566Validation lossGPT Medium00.50.5111.51.52233445566GPT XL
3
Figure 3: Validation loss for GPT XS/Small/Medium/XL across six optimizers. X-axis: token count (billions). 8-bit Muon optimizers are competitive with Muon, closely matching its performance.
Model AdamW-32 AdamW-8D Muon-32 M-8L/A-32 M-8D/A-32 M-8D
loss (% close) loss (% close) loss (% close)
XS $3.509$ $3.499$ 3.325 $3.359$ $1.02$ $3.362$ $1.10$ $3.364$ $1.16$
Small $3.406$ $3.489$ 3.192 $3.228$ $1.14$ $3.232$ $1.25$ $3.233$ $1.28$
Medium $3.269$ $3.624$ 3.030 $3.070$ $1.32$ $3.078$ $1.61$ $3.082$ $1.68$
XL $2.994$ $3.020$ $2.837$ $2.830$ 0.25-0.25 $2.822$ 0.53-0.53 2.803 1.20-1.20
Table 3: Final validation loss for pre-training across models and optimizers. M = Muon, A = Adam. All 8-bit Muon variants are within $\sim 1-2\%$ of Muon’s loss, and outperform all AdamW variants comprehensively. For XL, they match the performance of Muon.
5.2 Fine-tuning with 8-bit Muon
Architectures We use the Llama 3.2 3B base model (Dubey et al., 2024) for all our fine-tuning (SFT) experiments. It is empirically well-known that if a model is pre-trained with one optimizer, then empirically it is best to fine-tune it with the same optimizer Liu et al. (2025); Team et al. (2025). Our aim is to assess how closely quantized versions of AdamW and Muon can match their full precision counterparts, and not to achieve state-of-the-art results.
Datasets We use the tulu-3-sft-mixture dataset (Lambert et al., 2024) - an open dataset for post-training which targets a diverse set of skills such as reasoning and math. The dataset consists of close to 1 million training samples.
Training Details We use SFT on the Llama model using all the optimizers used for pre-training. For each optimizer, we use a random 10k split samples from the full Tulu-3 dataset to perform lightweight fine-tuning, since our aim is to compare optimizers in a fair setting. This data-constrained setup is ideal for our aim of comparing optimizers across varying training regimes. Since SFT of LLMs can have non-negligible variance, we conduct 5 runs per experiment.
We inherit all the optimizer setup from pre-training, except for learning rate and weight decay. Learning rate follows a linear warmup to $10^{-5}$ for the first 3% steps, followed by linear decay to 0. Weight decay is set to 0.01. We use a single NVIDIA H100 GPU for all experiments, with a batch size of 2 and gradient accumulation of 8, yielding a global batch size of 16 sequences.
Evaluation Criteria Our SFT evaluation is performed on the two benchmarks: HumanEval(pass@1) (Chen et al., 2021) for coding and GSM8K (Cobbe et al., 2021) for math.
5.2.1 Results
SFT results can be found in Table 4. Again, the 8-bit variants of Muon match the performance of the 32-bit version. Since Llama 3.2 3B is pre-trained with AdamW, there is no clear winner between AdamW and Muon, mirroring findings from other papers (Liu et al., 2025; Team et al., 2025). We expect Muon-based fine-tuning to be increasingly beneficial as more models pre-trained with Muon are released.
Model (GPT) Adam-32 Adam-8D Muon-32 M-8L/A-32 M-8D/A-32 M-8D
GS8K $28.30$ 0.61 $29.16$ 0.42 $28.05$ 0.425 $28.66$ 0.619 $28.253$ 0.57 $28.73$ 0.35
HumanEval $26.63$ 0.70 $27.64$ 1.86 $26.42$ 0.93 $27.64$ 0.35 $26.83$ 0.61 $26.95$ 1.17
Table 4: Llama 3.2 3B SFT Results after lightweight training. Standard deviation across 5 runs in shown in subscript 8-bit variants are competitive with their 32-bit versions.
5.3 Memory footprint
Table 5 compares the persistent HBM memory footprint of the optimizer state for variants, when profiled for the GPT Small, Medium and XL models, as well as Llama-3.2-3B Grattafiori et al. (2024) (Figure 1 summarizes the same info for XL and Llama). For the biggest models like XL and Llama, Muon-8D provides substantial relative savings of $\sim$74%, $\sim$86% and $\sim$44%, when compared to Muon-32, AdamW-32 and AdamW-8D respectively.
[…]
Table 5: Optimizer states memory (GB) across all model sizes tested (lower is better). Muon-8D reduces optimizer memory footprint for GPT XL and Llama substantially.
For smaller models, Muon-8D/Adam-32 has a larger optimizer state memory footprint than Adam-8D, a trend that inverts as model size increases. This is because the size of the embedding and lm-head matrices remains the same across model sizes because of a fixed vocabulary. For the XS model, these layers constitute up to 47% of model parameters and are optimized with 32-bit AdamW. The high memory cost for these layers offsets the savings achieved in the rest of the model.
@@ 6 Conclusions
In this paper, we introduced 8-bit Muon, a memory-efficient optimizer designed to address the problem of large memory footprints of LLMs. We build on the Muon optimizer and leverage blockwise quantization. One key finding is the robustness of Muon to types of quantization. Our results across pre-training and fine-tuning of large models show that our 8-bit Muon variants nearly matched the performance of the full-precision Muon. In terms of practical benefits, our method reduced the optimizer state memory by up to 86% compared to AdamW and 74% compared to full-precision Muon for models up 1.6B-3B in size. Future work could include quantization to even lower bits, as well as combination with techniques like low-rank matrices.
@@ A.1.2 Pre-training DETAILS
We train 4 GPT-style models on 2B to 4B tokens using 8 H100 GPUs with PyTorch Distributed DataParallel (DDP). A summary of the model architectures used is described in Table 6.
Model sizes up to 405M were trained on 2B tokens with a learning rate of $3\cdot 10^{-4}$ with a per-GPU batch size of 16 samples or 32768 tokens. We purposely use the same number of samples per optimizer step for all model sizes despite lower Model FLOPs Utilization to have a fair comparison between model sizes. The same logic is applied to the XL model, which is trained with a local batch size of 2 and 8 gradient accumulation steps for all optimizer variants. A summary of the training hyperparameters used for training is shown in Table 7
@@
XL 1.6B 1600 48 25 4
Table 6: Model sizes used for the pre-training task.
Model size LR Local BS Grad. Acc. Steps Global BS
XS $3\cdot 10^{-4}$ 32768 1 262144
Small $3\cdot 10^{-4}$ 32768 1 262144
Medium $3\cdot 10^{-4}$ 32768 1 262144
XL $2\cdot 10^{-4}$ 4096 8 262144
@@
Appendix B Muon algorithms
Algos 12 and 3 represent vanilla Muon (Jordan et al., 2024), Muon (Liu et al., 2025) and Quantized Muon. Note The algos use small-case vector notation.
Algorithm 1 Vanilla Muon (using $\bm{\theta},\mathbf{g},\mathbf{m}$; NS-orthogonalized momentum for 2D hidden-layer params)
1:Input: step size $\alpha$, momentum $\beta\in[0,1)$, NS steps $T$ (default 55), $\epsilon>0$
2:Initialize: $\bm{\theta}^{(0)}$, $\tilde{\mathbf{m}}^{(0)}=\mathbf{0}$
3:for $t=1,2,\dots$ do
4:  $\mathbf{g}^{(t)}\leftarrow\nabla f_{t}(\bm{\theta}^{(t-1)})$
5:  $\tilde{\mathbf{m}}^{(t)}\leftarrow\beta\tilde{\mathbf{m}}^{(t-1)}+(1-\beta)\,\mathbf{g}^{(t)}$ $\triangleright$ SGD-style momentum before orthogonalization
6:  for each 2D hidden-layer parameter block $\theta\subset\bm{\theta}$ do
7:   $\Delta\leftarrow\textsc{NS}(\tilde{\mathbf{m}}^{(t)}[\theta],\,T,\,\epsilon)$ $\triangleright$ Orthogonalize update
8:   $\theta\leftarrow\theta-\alpha\,\Delta$
9:  end for
10:end for
Algorithm 2 Muon (using $\bm{\theta},\mathbf{g},\mathbf{m}$; NS-orthogonalized momentum for 2D hidden-layer params)
1:Input: step size $\alpha$, momentum $\beta\in[0,1)$, weight decay $\lambda\geq 0$, NS steps $T$ (default 55), $\epsilon>0$
2:Initialize: $\bm{\theta}^{(0)}$, $\tilde{\mathbf{m}}^{(0)}=\mathbf{0}$
3:for $t=1,2,\dots$ do
4:  $\mathbf{g}^{(t)}\leftarrow\nabla f_{t}(\bm{\theta}^{(t-1)})$
5:  $\tilde{\mathbf{m}}^{(t)}\leftarrow\beta\tilde{\mathbf{m}}^{(t-1)}+(1-\beta)\,\mathbf{g}^{(t)}$ $\triangleright$ SGD-style momentum before orthogonalization
6:  for each 2D hidden-layer parameter block $\theta\subset\bm{\theta}$ (with dimensions $m,n$) do
7:   $\Delta\leftarrow\textsc{NS}(\tilde{\mathbf{m}}^{(t)}[\theta],\,T,\,\epsilon)$ $\triangleright$ Orthogonalize update
8:   $\theta\leftarrow(1-\alpha\lambda)\theta-0.2\alpha\sqrt{\max(m,n)}\,\Delta$
9:  end for
10:end for
Algorithm 3 Quantized Muon (using $\bm{\theta},\mathbf{g},\mathbf{m}$; NS-orthogonalized momentum for 2D hidden-layer params)
1:Input: step size $\alpha$, momentum $\beta\in[0,1)$, weight decay $\lambda\geq 0$, NS steps $T$, $\epsilon>0$, quantization params (mode, bits $B$)
2:Initialize: $\bm{\theta}^{(0)}$, quantized momentum state $(\mathbf{z}^{(0)},\mathbf{s}^{(0)})=(\mathbf{0},\mathbf{0})$
3:for $t=1,2,\dots$ do
4:  $\mathbf{g}^{(t)}\leftarrow\nabla f_{t}(\bm{\theta}^{(t-1)})$
5:  for each 2D hidden-layer parameter block $\theta\subset\bm{\theta}$ (with dimensions $m,n$) do
6:   $\tilde{\mathbf{m}}^{(t-1)}\leftarrow\textsc{Dequantize}(\mathbf{z}^{(t-1)}[\theta],\mathbf{s}^{(t-1)}[\theta])$ $\triangleright$ Dequantize saved momentum state
7:   $\mathbf{m}^{(t)}\leftarrow\beta\,\tilde{\mathbf{m}}^{(t-1)}+\mathbf{g}^{(t)}[\theta]$ $\triangleright$ Update momentum with current gradient
8:   $\mathbf{u}^{(t)}\leftarrow\textsc{NS}(\mathbf{m}^{(t)},T,\epsilon)$ $\triangleright$ Orthogonalize momentum for the update
9:   $\theta\leftarrow(1-\alpha\lambda)\theta-0.2\alpha\sqrt{\max(m,n)}\,\mathbf{u}^{(t)}$ $\triangleright$ Update parameters
10:   $(\mathbf{z}^{(t)}[\theta],\mathbf{s}^{(t)}[\theta])\leftarrow\textsc{Quantize}(\mathbf{m}^{(t)})$ $\triangleright$ Quantize and save new momentum state
11:  end for
12:end for
Appendix C Quantization Error Bounds
@@ C.1 Quantization Operator
C.2 Proofs
C.2.1 Proof of Theorem 1
@@
Summing over $i=1,\dots,d$ and using $\bm{\theta}^{(1)}-\tilde{\bm{\theta}}^{(1)}=\alpha\!\left(\frac{\mathbf{m}}{\sqrt{\mathbf{v}}+\epsilon}-\frac{Q(\mathbf{m})}{\sqrt{Q(\mathbf{v})}+\epsilon}\right)$,
$\mathbb{E}\|\bm{\theta}^{(1)}-\tilde{\bm{\theta}}^{(1)}\|_{2}^{2}$ $\geq\alpha^{2}d\nu\frac{g_{\infty}^{2}}{(256\epsilon)^{2}}.$
C.2.2 Proof of Theorem 2
@@
$\tilde{\bm{\theta}}^{(t+1)}-\bm{\theta}^{(t+1)}$ $=-\eta(\mathbf{g}^{(t)}+\rho Q(\mathbf{m}^{(t)}))+\eta(\mathbf{g}^{(t)}+\rho\mathbf{m}^{(t)})$
@@
$\|\tilde{\bm{\theta}}^{(t+1)}-\bm{\theta}^{(t+1)}\|_{2}^{2}$ $=\eta^{2}\rho^{2}\sum_{i=1}^{d}\left(Q(\mathbf{m}^{(t)})_{i}-\mathbf{m}_{i}^{(t)}\right)^{2}.$
@@
$\|\tilde{\bm{\theta}}^{(t+1)}-\bm{\theta}^{(t+1)}\|_{2}^{2}$ $\leq\eta^{2}\rho^{2}\sum_{i=1}^{d}\left(\|\mathbf{m}^{(t)}\|_{\infty}/127\right)^{2}$
@@ C.2.3 Proof of Theorem 3
C.3 Adam algorithm
For completeness, we reproduce the base Adam algorithm below (Algorithm 1 in Kingma (2014)).
Algorithm 4 Adam (using $\bm{\theta},\mathbf{g},\mathbf{m},\mathbf{v}$)
@@ C.3 Adam algorithm
2:Initialize: $\bm{\theta}^{(0)}$, $\tilde{\mathbf{m}}^{(0)}=\mathbf{0}$, $\tilde{\mathbf{v}}^{(0)}=\mathbf{0}$
@@
4:  $\mathbf{g}^{(t)}\leftarrow\nabla f_{t}(\bm{\theta}^{(t-1)})$
@@
9:  $\bm{\theta}^{(t)}\leftarrow\bm{\theta}^{(t-1)}-\alpha\,\mathbf{m}^{(t)}\big/(\sqrt{\mathbf{v}^{(t)}}+\epsilon)$
@@
C.4 Exact Muon optimizer

### v3

@@ 5.3 Linear Quantization of the Muon Optimizer
Consider a step of Muon with momentum (using the exact polar factor rather than the Newton–Schulz approximation) with and without 8-bit linear quantization, $Q$ (Definition 1), applied to the momentum. Let the layer weights and momentum state be $\mathbf{W}^{(t-1)}$ and $\mathbf{M}^{(t-1)}$, each with $d$ entries (see Appendix B.4 for full update formulas).
@@ 6 Conclusions
In this paper, we introduced 8-bit Muon, a memory-efficient variant of the Muon optimizer based on blockwise quantization of its optimizer state. We find that 8-bit Muon can match full-precision Muon in both validation loss and downstream benchmarks on Chinchilla-optimal GPT pretraining up to 2.7B parameters, while substantially reducing optimizer-state memory. In contrast to quantized AdamW, Muon remains compatible with simple linear quantization, providing a practical and easy-to-implement default. We complement these empirical results with a mechanistic and theoretical analysis that explains the differing behavior of AdamW, SGD with momentum, and Muon under quantization. Future work includes extending these methods to lower-bit quantization and combining them with complementary memory-saving techniques such as low rank momentum updates.
@@ A.1.2 Pre-training Details
We train five GPT2-style models using 8–32 B200 GPUs with PyTorch Distributed DataParallel (DDP). Table 7 summarizes the model architectures, and Table 8 summarizes the pretraining hyperparameters.
@@
Large 834M 1536 24 16 4
XL 1.4B 2048 24 16 4
XXL 2.7B 2560 32 32 4
Table 7: Model sizes used for pre-training. Parameter counts are slightly different than those reported by (Brown et al., $20$) as the full architectural details were not reported. As a result, subtle differences in architecture as vocabulary size, GQA vs MHA and others may impact final model size.
Model size LR Local BS Grad. Acc. Steps N. GPUs Global BS
Small $6\cdot 10^{-4}$ 65536 1 8 0.5M
Medium $3\cdot 10^{-4}$ 65536 1 8 0.5M
Large $2.5\cdot 10^{-4}$ 65536 1 8 0.5M
XL $2\cdot 10^{-4}$ 32768 1 32 1M
XXL $1.6\cdot 10^{-4}$ 32768 2 32 1M
@@
Appendix B Quantization Error Bounds
@@ B.1 Quantization Operator
B.2 Proofs
B.2.1 Proof of Theorem 1
@@
Summing over $i=1,\dots,d$ and using $\mathbf{w}^{(1)}-\tilde{\mathbf{w}}^{(1)}=\alpha\!\left(\frac{\mathbf{m}}{\sqrt{\mathbf{v}}+\epsilon}-\frac{Q(\mathbf{m})}{\sqrt{Q(\mathbf{v})}+\epsilon}\right)$,
$\mathbb{E}\|\mathbf{w}^{(1)}-\tilde{\mathbf{w}}^{(1)}\|_{2}^{2}$ $\geq\alpha^{2}d\nu\frac{g_{\infty}^{2}}{(256\epsilon)^{2}}.$
B.2.2 Proof of Theorem 2
@@
$\tilde{\mathbf{w}}^{(t+1)}-\mathbf{w}^{(t+1)}$ $=-\eta(\mathbf{g}^{(t)}+\rho Q(\mathbf{m}^{(t)}))+\eta(\mathbf{g}^{(t)}+\rho\mathbf{m}^{(t)})$
@@
$\|\tilde{\mathbf{w}}^{(t+1)}-\mathbf{w}^{(t+1)}\|_{2}^{2}$ $=\eta^{2}\rho^{2}\sum_{i=1}^{d}\left(Q(\mathbf{m}^{(t)})_{i}-\mathbf{m}_{i}^{(t)}\right)^{2}.$
@@
$\|\tilde{\mathbf{w}}^{(t+1)}-\mathbf{w}^{(t+1)}\|_{2}^{2}$ $\leq\eta^{2}\rho^{2}\sum_{i=1}^{d}\left(\|\mathbf{m}^{(t)}\|_{\infty}/127\right)^{2}$
@@ B.2.3 Proof of Theorem 3
B.3 Adam algorithm
For completeness, we reproduce the base Adam algorithm below (Algorithm 1 in (Kingma, 2014)).
Algorithm 1 Adam (using $\mathbf{w},\mathbf{g},\mathbf{m},\mathbf{v}$)
@@ B.3 Adam algorithm
2:Initialize: $\mathbf{w}^{(0)}$, $\tilde{\mathbf{m}}^{(0)}=\mathbf{0}$, $\tilde{\mathbf{v}}^{(0)}=\mathbf{0}$
@@
4:  $\mathbf{g}^{(t)}\leftarrow\nabla f_{t}(\mathbf{w}^{(t-1)})$
@@
9:  $\mathbf{w}^{(t)}\leftarrow\mathbf{w}^{(t-1)}-\alpha\,\mathbf{m}^{(t)}\big/(\sqrt{\mathbf{v}^{(t)}}+\epsilon)$
@@
B.4 Exact Muon optimizer
@@
B.5 Empirical Validation
In this section, we provide empirical validation for the theorems of Section 5.1 using a simple experimental setup. We train a two-layer fully-connected network with one hidden layer of width 256 and ReLU activation, mapping $28\times 28$ pixel inputs to 10 logits. Inputs are normalized with the standard MNIST mean and variance $(0.1307,0.3081)$, and we use the standard training split from torchvision with a batch size of 128 and cross-entropy loss.
Theorem 1 - We measure that approximately 29% of gradient entries in the loss gradient, $\mathbf{g}$, fall within the range $\|\mathbf{g}\|_{\infty}/60,\|\mathbf{g}\|_{\infty}/16]$ from a Kaiming Uniform initialization of the model weights. This indicates that the gradient coordinate assumption of Theorem 1 holds in practice, and hence provides a realistic explanation for the divergence observed after a single step of Adam.
Theorem 2 - We measure the quantization error $\|\tilde{\mathbf{w}}^{(t+1)}-\mathbf{w}^{(t+1)}\|_{2}^{2}$ as a proportion of the error bound in Theorem 2 for the SGD+M optimizer with learning rate $0.1$ and momentum $0.9$. For all $t$ over $1400$ training steps, the ratio of the true squared $\ell_{2}$ quantization error and the error bound stays between 0.06 and 0.085. By taking the square root of both sides in the error bound, we see that the true $\ell_{2}$-norm quantization error is accurately predicted by the theoretical bound up to a small constant factor.
Theorem 3 - We measure the quantization error $\|\widetilde{\mathbf{W}}^{(t)}-\mathbf{W}^{(t)}\|_{F}^{2}$ of the first layer weight matrix for the Muon optimizer with $\alpha=0.2$ and $\beta=0.95$. We observe that the ratio of the true squared Frobenius-norm quantization error to the bound of Theorem 3 stays between $2\times 10^{-6}$ and $1.4\times 10^{-5}$. The gap is explained by the fact that the proof of our theorem depends on the weak assumption that original and quantized matrices have full column rank $r$ with the $r$-th singular value of each matrix lower bounded by $s$. In fact, this condition is pessimistic, as the perturbation of the weight matrix is not aligned with the $r$-th singular vector. More complex assumptions on the spectrum of $\mathbf{W}$, paired with probabilistic assumptions on the quantization noise and a polar-factor perturbation analysis that accounts for additional spectral information (see Li and Sun (2003), for example), would produce a tighter bound. However, this analysis is outside the scope of this work.
B.6 Training and Validation Loss During Pretraining
Figure 3: Pretraining training-loss curves for GPT-2 Small, Medium, Large, XL, and XXL with three optimizers: Muon, Muon-8L, and Muon-8L/AdamW-8D. Curves show the mean over 5 random seeds; error bars (seed-to-seed variation) are present but visually negligible.
Figure 4: Pretraining validation-loss curves for GPT-2 Small, Medium, Large, XL, and XXL with three optimizers: Muon, Muon-8L, and Muon-8L/AdamW-8D. Curves show the mean over 5 random seeds; error bars (seed-to-seed variation) are present but visually negligible.
Figure 5: Relative pairwise validation-loss differences relative to the Muon baseline during pretraining for GPT-2 Small, Medium, Large, XL, and XXL.